calculate the wavelength associated with motion of a 3.2*10^(-22)g particle moving with a speed of half the speed of light
Answer #1

According to De. Broglie. th mv it is the wavelength his Plancks constant = 6.626 X 10 8 J- see m is the mass. V is the velo

V= 1.5*108m/s

Wavelength=(6.626*10-34 J. s) /(3.2*10-25 Kg*1.5*108m.s-1)

Wavelength=1.3804167*10-17 m